# Video-Skill-CoT: Skill-based Chain-of-Thoughts for Domain-Adaptive Video Reasoning

### v1

@@ C.2 Selected skills over different video datasets
In Figs. 3 and 4, we present statistics on the selected top 3 assigned skills for each task in VSI-Bench (presented in Sec. 3.2). As shown in the results, object identification skills are commonly used across tasks. However, each task also requires domain-specific skills. For instance, the Room Size Estimation task necessitates skills such as “Determining room boundaries using structural elements like walls and floors.
C.3 Inference rationale comparison
In Fig. 6, we compare the inference-time rationales generated by LLaVA-Video trained with (a) regular CoT and (b) our proposed skill-based CoT. During inference, we prompt each model with: “Explain the rationale to answer the question and answer the question.” As shown in (a), the model trained with regular CoT produces an incorrect reasoning process, ultimately leading to a wrong answer. In contrast, Video-SkoT successfully generates temporally grounded and precise rationales that more effectively support accurate answer generation.
C.4 Additional comparison with regular CoT
In Fig. 5, we provide additional comparison with regular CoT and ours.
@@ Appendix D License
Figure 7: Skill description from different domain datasets. We visualize the skill descriptions for each dataset after performing skill extraction and clustering. (Sec. 3.2)

### v2

@@ C.2 Selected skills over different video datasets
In Figs.˜4 and 5, we present statistics on the selected top 3 assigned skills for each task in VSI-Bench (presented in Sec.˜3.2). As shown in the results, object identification skills are commonly used across tasks. However, each task also requires domain-specific skills. For instance, the Room Size Estimation task necessitates skills such as “Determining room boundaries using structural elements like walls and floors.
C.3 Additional comparison with regular CoT
In Fig.˜6, we provide additional comparison with regular CoT and ours.
@@ Appendix D License
Figure 7: Skill description from different domain datasets. We visualize the skill descriptions for each dataset after performing skill extraction and clustering. (Sec.˜3.2)